Draw the graph of the function f(x) = 2x + 3 from the set of real numbers to the set
of real numbers.
1
Expert's answer
2020-09-07T16:14:11-0400

Since "f(x)" is a linear fucnction, we can obtain the graph of the line "f(x) = 2x +3" with calculating 2 points which belong to it, we should find 2 points on this line. Let "x = 0", then "f(x) = 2\cdot 0 + 3 = 3", and we obtain point "(0,3) = A" , and for "x = 1:" "f(x) = 2 \cdot 1 + 3 = 5," and we obtain point "(1, 5) = B" . Thus, we have the line through these points, which the graph of "f(x) = 2x +3"